OpenStudy (luigi0210):

Find the anti derivative of f''(x)= x^-2 x>0, f(1)=0, f(2)=0

sam (.sam.):

\[\int\limits \frac{1}{x^2}dx\] Do you know how to do this?

sam (.sam.):

\[\int\limits x^{-1}dx \\ \\ \] Apply \[\int\limits x^n dx=\frac{1}{n+1}x^{n+1}+c\]

sam (.sam.):

\[\int\limits\limits x^{-2}dx\] Apply \[\int\limits\limits x^n dx=\frac{1}{n+1}x^{n+1}+c\] So \[\int\limits \frac{1}{x^2} \, dx=-\frac{1}{x}+c\]